a firm is considering adopting a new technology to improve its production process. the implementation cost would be $300,000. The initial annual operating cost of $40,000 will increase by $5,000 per year after the first year. The new technology would produce yearly savings of $100,000. The time span before the technology becomes obsolete and needs to be replaced is estimated in 10 years. at the time of replacements, the salvage value of the obsolete equipment is estimated to be $ 50,000.
What rate of return would the firm make on this investment?
Answer #1

Answer:

In order to find the rate of return we will equate the pw to zero.

pw = implementation cost + aoc(p/a,i,n) + increase in aoc(p/g,i,n) + savings(p/a,i,n) + salvage value(p/f,i,n)

0 = -300,000 - 40,000(p/a,i,10) - 5,000(p/g,i,10) + 100,000(p/a,i,10) + 50,000(p/f,i,10)

300,000 + 40,000(p/a,i,10) + 5,000(p/g,i,10) = 100,000(p/a,i,10) + 50,000(p/f,i,10)

solving via trial and error we get that i is between 7% and 8% as at 7% , the pw is above zero and at 8% it is below zero and solving further we get that i is 7.66%

so the rate of return is 7.66%